Georgia Craig Below is the end of year trial balance for 2014 and 2015. All amounts are in their normal balance You are also provided 2015 CRI and CP All sales are credit All purchases are for inventory and are on credit There is no other income or expenses GP is 30% Calculate the missing amounts Owners Equity 2014 For 2015 Cash at Bank Inventory Sales COGS 2014 50000 5000 8000 10000 2015 Cash at Bank Accounts Receivable Inventory Accounts Payable Owners Equity Sales COGS 35000 75000 DR Cash 170000 CR Cash Receipts Journal 2015 Accts Receivable 170000
DR Cash 170000 CR Accts Receivable 170000 Cash Receipts Journal 2015 Cash 85000 Cash Payments Journal 2015 Accts Pay 85000
Answer #1
Owner's equity 2014
Assets=Liabilities+Equity
Equity=Assets-Liabilities
Assets:
$
Cash at bank 50000
Accounts receivable 5000
Inventory 8000
Total assets 63000
Liabilities=Accounts payable= $ 10000
Equity=Owner's equity (Due to lack of information)
Owner's equity=63000-10000=$ 53000
2015
Cash at bank:
$
Beginnig balance 50000
Add: Cash receipts 170000
220000
Less: Cash payments 85000
Ending balance 135000
Cash at bank=$ 135000
Sales
All sales are on credit
Hence, Balancing figure in accounts receivable will be sales
$
Ending balance 35000
Add: Cash received 170000
205000
Less: Beginning balance 5000
Sales 200000
Sales= $ 200000
COGS
GP is 30%
Hence, COGS=100-30=70%
COGS=Sales*70%=200000*70%=$ 140000
Inventory:
COGS=Beginning inventory+Purchases-Ending inventory
Ending inventory=Beginning inventory+Purchases-COGS
All purchases are for inventory and on credit.
Hence, Balancing figure in accounts payable will be purchase
$
Ending balance 75000
Add: Cash paid 85000
160000
Less: Beginning balance 10000
Purchase 150000
Ending inventory=8000+150000-140000=$ 18000